OpenStudy (anonymous):

Help me for a medal :D? Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 |5x − 6| = −41 Equation 2 |7x + 13| = 27 Equation 1 has more solutions than equation 2. Equation 2 has more solutions than Equation 1. I've done this twice, and I got it wrong so I removed the possible answers that I knew were wrong.

OpenStudy (anonymous):

absolute value problems ... 5x -6 = -41 5x (-6+6) = -41+6 then you have 5x = a negative number and to solve with division; you will have your answer for equation 2 use the same process: 7x +13 =27 this goes to 7x = 27-13 and with the absolute value in place and your answer positive, you will arrive at a pair of answer possible. Hope this helps.